OpenStudy (anonymous):

Find the margin of error for a survey that has the given sample size. Round the answer to the nearest tenth of a percent. 1. 225 2. 100 3. 625 4. 3600 5. 4200 How do i do this, please helpp.

OpenStudy (rogue):

According to this great video I found, you'd need more information to do this problem. In addition to the sample size, you need the confidence coefficient and standard deviation. http://www.youtube.com/watch?v=Mfia4nbh-zU

OpenStudy (anonymous):

it will be 2nd one

OpenStudy (anonymous):

My professor did some of these problems today & this was all the information he had & he solved it.

OpenStudy (rogue):

Hmm, then assume standard deviation and confidence level to be 1? :/ If so, \[margin \space of \space error = \frac {1}{\sqrt {sample \space size}}\]

OpenStudy (rogue):

Normally, it is \[margin \space of \space error = z \times \frac {\sigma}{\sqrt n}\]